OpenStudy (anonymous):

CAN SOMEONE PLEASE HELP ME WITH THIS ONE QUESTIONS??? Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution. x plus a over ax = b over x Part 1. Show all work to solve for x in the equation and check the solution.

OpenStudy (anonymous):

@satellite73

OpenStudy (emmigrace222):

hmm let me look do you have anymore information you could give me?

OpenStudy (anonymous):

\[\frac{x+a}{ax}=\frac{b}{x}\]?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Not reeally, all i know that anextraneous solution is a solution of a simplified version of an equation that does not satisfy the original equation

OpenStudy (anonymous):

what are you supposed to do? find \(a\) and \(b\) with no solution ?

OpenStudy (anonymous):

THey want me to plug in numbers in a and b

OpenStudy (anonymous):

that has an extraneous solution

OpenStudy (emmigrace222):

@satellite73 is awesome and super smart :) ill leave y'all alone :)

OpenStudy (anonymous):

you could make them both 1 that wouldn't have a solution but i am not sure what it is really askign

OpenStudy (anonymous):

\[\frac{x+1}{x}=\frac{1}{x}\] has no solution

OpenStudy (anonymous):

my teacher told me to plug in values into the equation and find extraneous solution but im confused

OpenStudy (anonymous):

being simple minded i made \(a=b=1\)

OpenStudy (anonymous):

You would have to cross multiply after?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Okay, one second, im going to write it down on paper

OpenStudy (anonymous):

so would it be x(x+1)= x1?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

What about when you solve it? would it be 0?

OpenStudy (anonymous):

yes it would

OpenStudy (anonymous):

can you show me the steps though?

OpenStudy (anonymous):

you would ge t \[x^2+x=x\\ x^2=0\\ x=0\] but that is not a solution

OpenStudy (anonymous):

oh ok, thanks

OpenStudy (anonymous):

and \(0\) in not a solution to the original equation because you cannot have a zero in the denominator

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

So is it okay, if i use these steps to write it down?

OpenStudy (anonymous):

sure why not?

OpenStudy (anonymous):

Okay, thank you very much for the help